# Exploring Model Kinship for Merging Large Language Models

### v2

@@ Abstract
Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs). The open-source community has driven model evolution by iteratively merging existing models. However, a principled understanding of the expected gains and underlying factors in model merging remains lacking. In this work, we examine model evolution through continual merging, analogous to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs. With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models. Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can yield better performance on benchmark datasets. Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
@@
Yedi Hu, Yunzhi Yao, Shumin Deng , Huajun Chen, Ningyu Zhangthanks: Corresponding Author. Zhejiang University   National University of Singapore, NUS-NCS Joint Lab, Singapore {zhangningyu}@zju.edu.cn
@@ 1 Introduction
Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023). However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang and Yang, 2022; Lu et al., 2024; Liu et al., 2024).
@@
Figure 1: An intuitive comparison between wheat evolution and model evolution. An interesting parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b). In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits. Similarly, in model merging, the merged model inherits parameters or weights from the contributing models. Part c demonstrates the iterative execution of model evolution. Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration. Notably, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
@@
Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multitask objectives by integrating multiple expert models. Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
@@
To date, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023). Many of these models are created through a biologically inspired evolutionary process that involves iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)). Despite these successes, the current approach faces critical limitations. Progress often relies on trial and error and extensive human expertise, with little formal guidance or standardized procedure. To address this problem, we propose a strategic model evolution framework (Figure 1(c)) that leverages explicit strategies to guide the direction of model evolution toward improved performance. We show that even a simple greedy strategy can outperform baseline merging approaches.
@@
However, in the late stages of both the community experiments and the greedy strategy evolution, achieving further gains in generalization becomes increasingly difficult. To explore possible solutions, we introduce ’model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process. Model kinship is designed to quantify the degree of similarity or relationship between models throughout the iterative merging process. By providing a principled way to measure these relationships, model kinship offers crucial insights for refining merging strategies.
@@
We conduct a comprehensive analysis of model merging experiments based on model kinship. We observe that the community model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau. Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of the potential effectiveness of a merge. These findings inspire two main insights: (1) high-kinship merges can lead to performance stagnation, akin to inbreeding; (2) low-kinship merges carry greater risk but may yield larger gains and facilitate escape from local optima.
@@
Inspired by this, we propose a new continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship.  [^2] https://github.com/zjunlp/ModelKinship.. Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping mitigate degradation and avoid local optima during model evolution. Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process.
In general, this paper makes four key contributions:
@@
    Continual Model Merging as a Feasible Framework for Model Evolution: We propose continual model merging as a viable framework for evolving large language models. Through strategically guided merging across iterations, this approach yields consistent improvements in generalization and task performance.
@@
    Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging. Our findings highlight the dynamics of multitask performance improvement and stagnation during evolution. In addition, we propose a preliminary explanation of the underlying mechanisms using model kinship.
@@
    Practical Model Merging Strategies using Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
@@ 2.1 Model Merging: Fundamentals
2.2 Continual Model Merging: Benefits and Challenges
Parameter averaging methods enable the merged model to retain the same architecture and parameter size as the original models, allowing for reuse in future merging processes. Leveraging this property, the community has progressively enhanced models through repeated applications of model merging, a process we refer to as “Model Evolution”. Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that model evolution can yield highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
However, current model evolution practices typically involve random merging by multiple contributors, leading to high computational costs and unstable behavior that limit its feasibility for industrial applications.
3 Strategic Model Evolution
In this section, we demonstrate how strategic model evolution via continual merging can achieve better generalization across tasks.
3.1 Method
We first define an iterative process for model evolution, where models are incrementally merged to produce a more generalized model. The direction of model evolution is guided by a principled strategy that governs the model selection. At each iteration $t$, a selection strategy $\mathcal{S}$ is applied to the current pool of models $\mathcal{P}_{t}=\{M_{t}^{(1)},M_{t}^{(2)},\ldots,M_{t}^{(k)}\}$, producing a subset of candidate models $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$. These selected models are then combined using a merge method $\mathcal{M}$, such as weighted averaging or linear interpolation in the weight space, resulting in the next-generation model $M_{t+1}$. This iterative process continues until a predefined end strategy $\mathcal{E}$ is matched. Formally, the process is defined as:
$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{subject to}\quad\neg\mathcal{E}(M_{t+1})$ (2)
@@ 3.1 Method
Ties 62.76 78.61 0.11 47.16
Linear 56.37 78.08 68.54 67.66
Continual Merge (ord1) 47.15 76.24 53.15 58.84
Continual Merge (ord2) 61.01 79.56 63.76 68.11
Continual Merge (ord3) 49.80 78.53 55.72 61.35
Strategic Model Evolution 50.94 80.11 75.13 68.72
Table 1: Performance Comparison across different merging methods.
@@ 3.2 Setup
Baseline Method.
We adopt two commonly used approaches for merging multiple models as baselines. The first is multi-model merging, where all models are merged simultaneously. The second is sequential continual merging, where three models are merged in sequence using a pairwise merging strategy. For these baselines, we apply TIES (Yadav et al., 2023) and Linear Merge for the multi-model merging case and SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for the pairwise merging in the sequential strategy.
Top $k$ Greedy Merging.
Our method applies continual model merging with greedy strategy, the Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1 in Appendix C). Each single merging step in the continual merging experiments is performed using SLERP.
@@
We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
@@
Evaluation is performed using three task-specific benchmark datasets: Winogrande, GSM8k, and TruthfulQA. These benchmarks demonstrate the distinct strengths of the three selected fine-tuned models. Further details on the tasks are provided in Appendix A.3.
@@ 3.3 Results
As shown in Table 1, Strategic Model Evolution can yield better generalization when combining multiple tasks. However, compared to simple sequential merging, the improvements are limited due to the simplicity of the greedy strategy. This observation raises an important question regarding optimal merging strategies:
Problem: How can we design more principled and effective strategies for model merging to better facilitate model evolution?
@@ 4 Preliminary Analysis of Model Kinship
To explore efficient strategies for model evolution, we propose a novel metric, model kinship, which captures task-related differences between models and may help predict merging outcomes. We then conduct a preliminary analysis of model kinship in community merging experiments on LLMs to evaluate its potential for informing and enhancing model evolution strategies.
@@ 4.1 Model Kinship
Drawing inspiration from the parallel between artificial selection and model evolution (as detailed in Appendix G), we hypothesize that a concept analogous to kinship, the genetic relatedness studied in evolutionary biology (Thompson, 1985), can also apply to model merging. Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between models. This analogy suggests that just as genetic kinship influences breeding outcomes, model kinship may similarly impact the effectiveness of merging strategies in enhancing generalization performance.
@@
Model kinship $r$ is designed to capture the similarity of task capabilities between models. In this paper, we explore multiple potential metrics for evaluating similarity. For the calculation, $sim(\cdot,\cdot)$ denotes the similarity metric function used. Considering two cases merging of 2 models and merging of $n$ models, we formally define model kinship $r$ as:
@@ 4.2 Evaluation Metrics
Let $T$ be the set of tasks in the task group, where $T=\{T_{1},T_{2},\ldots,T_{n}\}$. Each task $T_{i}$ in the set $T$ is associated with a performance measure $P_{i}$ for the LLM. For a multitask objective, the Average Task Performance (Avg.) $\bar{P}$ is calculated using the equation:
@@ 4.3 Analysis of Model Kinship: Correlation and Evolution Dynamics
In this section, we analyze model kinship from two perspectives: (1) its correlation with performance gain across a broad range of open-sourced LLM merges, and (2) its dynamic along specific model evolution paths. These analyses aim to clarify the relationship between model kinship and multitask capability improvements, as well as to identify phases of merge effectiveness.
@@ 4.3.1 Correlation Between Model Kinship and Performance Gain
We begin by exploring the potential relationship between merge gain and model kinship using three similarity metrics: Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED). The models used in this analysis are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
@@ 4.3.2 Model Kinship in Evolution Paths
Figure 5: Left (a): The comparison of task performance improvement across merging generations. The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities. Right (b): The partial model family tree from the controlled experiments. The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
@@
  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multitask balance.
@@
Figure 4 shows how the model kinship and normalized average performance change over the course of the merging process. Both metrics display a similar two-stage trend: they increase during the Improving Stage and level off during the Saturation Stage. This parallel trajectory suggests a potential link between model kinship and performance gains, indicating that improvements in generalization may coincide with and possibly depend on increases in model kinship.
@@
Additionally, to generalize the findings from the Evolution Paths, we analyze how model kinship evolves across different stages of the merging process, thereby broadening the scope from individual paths to the full evolution picture. (detailed in Appendix D.1). The results show that the best performing models exhibit high mutual kinship, which may lead to a stagnation stage in the merging process.

### v3

@@ Abstract
Model merging has emerged as a key technique for enhancing the capabilities and efficiency of Large Language Models (LLMs). The open-source community has driven model evolution by iteratively merging existing models, yet a principled understanding of the gains and underlying factors in model merging remains limited. In this work, we study model evolution through iterative merging, drawing an analogy to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs. Through comprehensive empirical analysis, we show that model kinship is closely linked to the performance improvements achieved by merging, providing a useful criterion for selecting candidate models. Building on this insight, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can improve benchmark performance. Specifically, we discover that incorporating model kinship as a guiding criterion enables continuous merging while mitigating performance degradation caused by local optima, thereby facilitating more effective model evolution [^1] Code is available at https://github.com/zjunlp/ModelKinship..
@@
Yedi Hu, Yunzhi Yao, Ningyu Zhang, Huajun Chen, Shumin Dengthanks: Corresponding Author. Zhejiang University   National University of Singapore, NUS-NCS Joint Lab {231sm,zhangningyu}@zju.edu.cnshumin@nus.edu.sg
@@ 1 Introduction
Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, and has proven particularly effective for Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023). However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives increasing demands for multi-task learning solutions (Zhang and Yang, 2022; Yu et al., 2024a; Lu et al., 2024; Liu et al., 2024).
@@
Figure 1: An intuitive comparison between wheat evolution and model evolution. A parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b). In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits. Similarly, in model merging, the merged model inherits parameters or weights from the contributing models. Part c demonstrates the iterative execution of model evolution. Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration. Notably, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
@@
Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multi-task objectives by integrating multiple expert models. Furthermore, advances in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) have lowered the technical barrier, allowing users with limited expertise to easily conduct merging experiments, thus leading to an evolution of LLMs for the community.
@@
At present, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023). Many of these models are created through a biologically inspired evolutionary process involving iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)). Despite these successes, current merging practice faces critical limitations. Progress often relies on continuous trial and extensive human expertise, with little formal guidance or standardized procedure. To address this problem, we propose an iterative model merging framework (Figure 1(c)), leveraging explicit strategies to guide the direction of model evolution toward improved performance. We show that even a simple greedy strategy can outperform baseline merging approaches.
@@
However, in the later stages of both community-driven model merging experiments and greedy evolution strategies, achieving additional gains in multi-task capability becomes increasingly challenging. To explore possible solutions, we introduce ‘model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process. Model kinship is designed to quantify the degree of similarity or relationship between models throughout the iterative merging process. By offering a principled framework for measuring these relationships, model kinship provides valuable insights that can refine merging strategies for model evolution.
@@
We conduct a comprehensive analysis of model merging experiments based on model kinship. We observe that the model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau. Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of potential effectiveness for model merging. These findings inspire two main insights: (1) high-kinship merges can lead to performance stagnation, akin to inbreeding; (2) low-kinship merges carry greater risk but may yield larger gains and facilitate escape from local optima.
@@
Inspired by this, we propose a novel continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship. Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping mitigate degradation and avoid local optima during model evolution. Furthermore, model kinship also proves useful as an early stopping criterion, improving efficiency of the merging process.
In general, this paper mainly contains four key contributions:
@@
    Iterative Model Merging as a Framework for Model Evolution: We propose continual model merging as a viable framework for evolving LLMs. Through strategically guided merging across iterations, this approach yields consistent improvements in generalization and task performance.
@@
    Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging. Our findings highlight the dynamics of multi-task performance improvement and stagnation during evolution. In addition, we propose a preliminary explanation of the underlying mechanisms using model kinship.
@@
    Practical Model Merging Strategies with Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
@@ 2.1 Model Merging: Fundamentals
2.2 Model Evolution: Benefits and Challenges
Parameter averaging methods allow merged models to retain the same architecture and parameter size as their original components, enabling reuse in subsequent merging steps. Leveraging this property, the open-source community has progressively enhanced model performance through repeated merging, leading to a “Model Evolution”. Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that this iterative model merging process can produce highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
Despite these advances, current community practices largely rely on random merging by multiple contributors, which results in high computational costs and unstable behavior, limiting their practicality for industrial applications.
3 Iterative Model Merging and Strategy Boost
In this section, we present controlled experiments to demonstrate that iterative model merging, the process of repeatedly combining models, can help stabilize improvements in multi-task capability. Moreover, we demonstrate that incorporating a selection strategy, such as Top-$k$ Greedy, yields more substantial performance improvements.
3.1 Iterative Model Merging Framework
We formally define iterative model merging as a modular framework that evolves a population of models through repeated merging. It is governed by three components: a selection strategy $\mathcal{S}$, which selects candidate models from the pool $\mathcal{P}_{t}$; a merging operator $\mathcal{M}$, such as SLERP or weighted averaging, to combine the selected models; and an stopping criterion $\mathcal{E}$, which determines when to stop the process.
At each generation $t$, a subset $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$ is selected using $\mathcal{S}$, merged via $\mathcal{M}$, and the resulting model $M_{t+1}$ is added back to the pool. The process continues until the stopping criterion $\mathcal{E}$ is met:
$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{until}\quad\mathcal{E}(M_{t+1})$ (2)
@@ 3.1 Iterative Model Merging Framework
Ties (Density=0.5, Weight=0.3) 62.76 79.56 8.79 50.37
Dare Ties (Density=0.5, Weight=0.3) 59.36 79.08 65.73 68.05
Linear (Weight=0.3) 56.37 78.08 68.54 67.66
Sequential SLERP Merge (ord1) 47.15 76.24 53.15 58.84
Sequential SLERP Merge (ord2) 61.01 79.56 63.76 68.11
Sequential SLERP Merge (ord3) 49.80 78.53 55.72 61.35
Random Merging (k=3) 54.32 78.53 72.81 68.55
Top k Greedy Merging (k=3) 50.94 80.11 75.13 68.72
Table 1: Performance Comparison across multi-model merging, sequential model merging with different SLERP orders, and iterative model merging with strategy.
@@ 3.2 Setup
Baseline Methods.
We consider two types of baselines: (1) Multi-model merging methods that support merging all models at once, such as TIES (Yadav et al., 2023), Dare-TIES (Yu et al., 2024b), and Linear; and (2) Sequential merging, where models are merged pairwise using SLERP (Shoemake, 1985).
Top-$k$ Greedy Merging.
Our approach applies iterative model merging using a Top-$k$ greedy selection strategy on $n$ LLMs (see Algorithm 1). Each merge step uses SLERP. We also include a random merge baseline (Appendix C.2).
@@
We use three HuggingFace LLMs based on Mistral-7B: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
@@
Models are evaluated on Winogrande, GSM8k, and TruthfulQA, which highlight their task-specific strengths. Dataset details are in Appendix A.3.
@@ 3.3 Results
As shown in Table 1, iterative model merging can yield better generalization when combining multiple tasks. In particular, both random and Top-$k$ greedy iterative merging outperform one-step baselines, demonstrating the effectiveness of continual merging. While single-step merging methods can sometimes achieve strong results (e.g., Linear or Dare Ties), their performance is often unstable and sensitive to hyperparameter. Ties, for instance, fails drastically on GSM8k. Sequential SLERP merging shows similar limitations, as its performance varies significantly depending on the merge order. In contrast, iterative merging strategies are more stable, consistently yielding robust performance across tasks.
However, these methods offer only marginal improvements. This raises a key question: How can we further enhance performance beyond the current limits of naive or greedy strategies?
@@ 4 Preliminary Analysis of Model Kinship
To address this limitation, we investigate whether structural signals beyond raw performance can guide model selection. We introduce model kinship, a metric capturing parameter-space similarity, and conduct a preliminary analysis to examine its correlation with merge gains in community-merged LLMs. This helps assess the potential of kinship-aware strategies in guiding continual merging and avoiding local optima.
@@ 4.1 Model Kinship
Drawing inspiration from the parallel between artificial selection and model evolution (as detailed in Appendix G), we hypothesize that a concept analogous to kinship, the genetic relatedness studied in evolutionary biology (Thompson, 1985), can also apply to model merging. Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between the merge candidates. This analogy suggests that, as genetic kinship affects breeding outcomes, model kinship similarly influences the effectiveness of merging strategies in enhancing multi-task performance.
@@
Model kinship $r$ is designed to capture the similarity of task capabilities between models. In this paper, we explore multiple potential metrics for evaluating similarity. For the calculation, $sim(\cdot,\cdot)$ denotes the similarity metric function used. Considering two cases merging of 2 models and merging of $n$ models, we define model kinship $r$ as:
@@ 4.2 Evaluation Metrics
Let $T$ be the set of tasks in the task group, where $T=\{T_{1},T_{2},\ldots,T_{n}\}$. Each task $T_{i}$ in the set $T$ is associated with a performance measure $P_{i}$ for the LLM. For a multi-task objective, the Average Task Performance (Avg.) $\bar{P}$ is calculated by:
@@ 4.3 Analysis of Model Kinship: Correlation and Evolution Dynamics
In this section, we analyze model kinship from two perspectives: (1) its correlation with performance gain across a broad range of open-sourced LLM merges, and (2) its dynamic along specific model evolution paths. These analyses aim to clarify the relationship between model kinship and multi-task capability improvements, as well as to identify phases of merge effectiveness.
@@ 4.3.1 Correlation Between Model Kinship and Performance Gain
We begin by exploring the potential relationship between merge gain and model kinship using three similarity metrics: Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED). The models used are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
@@ 4.3.2 Model Kinship in Evolution Paths
Figure 5: Left (a): The comparison of task performance improvement across merging generations. The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multi-task capabilities. Right (b): The partial model family tree from controlled experiments. The red arrow shows critical changes between experiment 1 and 2 in the evolution path.
@@
  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multi-task balance.
@@
Figure 4 shows how the model kinship and normalized average performance change over the course of the merging process. Both metrics exhibit a consistent two-phase trend: an Improving Stage, characterized by steady increases, followed by a Saturation Stage, where growth plateaus. This parallel progression highlights a potential correlation between model kinship and performance gains, indicating that enhancements in generalization are not only concurrent with but may also be facilitated by increases in model kinship.
@@
Greedy Strategy + Model Kinship
[…]
MetaMath 63.72 / / / MetaMath 63.72 / / /
Instruct 61.82 / / / Instruct 61.82 / / /
Open-chat 66.28 / / / Open-chat 66.28 / / /
model-1-1 62.17 -0.6 -4.11 0.01 model-1-1 62.17 -0.6 -4.11 0.01
model-1-2 64.02 -0.03 -2.26 -0.02 model-1-2 64.02 -0.03 -2.26 -0.02
model-1-3 66.84 +1.84 +0.56 0.05 model-1-3 66.84 +1.84 +0.56 0.05
model-2-1 68.72 +2.16 +1.88 0.93 model-2-1 68.72 +2.16 +1.88 0.93
model-2-2 61.47 -3.96 -7.25 0.57 model-2-2 61.47 -3.96 -7.25 0.57
model-2-3 61.32 -3.83 -7.40 0.58 model-2-3 61.32 -3.83 -7.40 0.58
model-3-1 68.59 +1.09 -0.13 0.95 model-3-2 67.74 +1.09 -0.98 0.93
model-3-2 67.74 -0.04 -0.98 0.93 model-3-3 69.06 +0.74 +0.34 0.24
- - - - model-3-4 68.61 +1.13 -0.11 0.32
model-4-1 68.51 -0.14 -0.08 0.98 model-4-4 68.75 -0.14 -0.31 0.54
model-4-2 68.04 -0.19 -0.67 0.98 model-4-5 68.39 -0.27 -0.36 0.66
model-4-3 68.53 +0.37 -0.06 0.94 model-4-6 69.03 +0.15 +0.42 0.52
- - - - model-5-1 69.13 +0.04 +0.10 0.65
- - - - model-5-2 68.98 +0.07 -0.15 0.65
- - - - model-5-3 68.63 -0.37 -0.50 0.98
To move beyond individual Evolution Paths, we further investigate how model kinship develops across different stages of the merging process, thereby extending our analysis to the broader evolutionary landscape (refer to Appendix D.1). Our findings reveal that the highest-performing models maintain strong mutual kinship. However, this close relatedness can also induce a stagnation phase, where the lack of diversity limits further performance improvements.